OpenStudy (anonymous):

I know that x+3=5x-4 (x=7/4) is equal to 10x+6=2x-8 (x=14/8). Is there way to tell if they have a the same solution without solving the equations?

OpenStudy (anonymous):

I wouldn't think so. Solving it seems to be better since you will know for sure what the answer is.

OpenStudy (anonymous):

Unless I'm wrong of course

OpenStudy (texaschic101):

10x + 6 = 2x - 8 --- reduce...divide everything by 2 5x + 3 = x - 4 and 5x + 3 = x - 4 IS NOT the same as x + 3 = 5x - 4